# MLOps with Microservices:A Case Study on the Maritime Domain

### v1

@@ Abstract
This case study describes challenges and lessons learned on building Ocean Guard: a Machine Learning–Enabled System (MLES) for anomaly detection in the maritime domain. First, the paper presents the system’s specification, and architecture. Ocean Guard was designed with a microservices’ architecture to enable multiple teams to work on the project in parallel. Then, the paper discusses how the developers adapted contract-based design to MLOps for achieving that goal. As a MLES, Ocean Guard employs code, model, and data contracts to establish guidelines between its services. This case study hopes to inspire software engineers, machine learning engineers, and data scientists to leverage similar approaches for their systems.
@@ 1 Introduction
Since 2015, the microservices architectural style have emerged as a promising approach to address challenges of modularity, scalability, and resilience by decomposing complex systems into smaller, independently deployable services with well-defined interfaces [14, 16]. This approach can be particularly suited to the maritime domain’s nature, where specialized components can be designed to handle different maritime data sources and analytical needs.
@@
This case study describes the development of Ocean Guard: a ML-Enabled System (MLES) for anomaly detection in the maritime domain. Section 2 details the functional and non-functional requirements considered for the system. Section 3 gives an overview of the system architecture, designed with a microservices architecture for MLOps. Section 4 zooms in in the Ocean Guard API, designed with a hexagonal architecture with Domain Driven Design (DDD). Section 5 shows the different teams involved in the development of the tool. Section 6 describes how the developers adapted contract-based design to MLES for developing different components in parallel. Section 7 lists the main challenges during the development of the Ocean Guard tool, according to the experience of the authors. Section 8 then summarizes the key lessons learned, related to applying development techniques from microservices on ML-enabled system development. Finally, section 9 lists the main contributions of paper and future work.
@@ 2 System Specification
This section describes the specification of the Ocean Guard tool, which was defined from conversations with stakeholders at the beginning of the MARIT-D project. The specification is divided into two parts: functional and non-functional requirements [15].
@@ 2.1 Functional Requirements
Functional Requirements, document the functionalities that directly affect the user of a system [15]. There are two actors related to the Ocean Guard tool. The following descriptions detail the main functionalities associated with them.
@@
    see geolocations of marine objects in a map,
@@ 2.2 Non-Functional Requirements
Non-Functional Requirements, also known as Architectural Requirements, describe a set of nonstructural attributes that are important for the success of a system [15]. There are four main non-functional requirements proposed for the Ocean Guard tool. The following descriptions detail the main characteristics the system should have to attend its users.
Compliance.
The Ocean Guard tool should be compliant with the latest European laws, including the General Data Protection Regulation (GDPR), and the Artificial Intelligence Act (AI Act).
@@
Compatibility.
@@ 3 System Architecture
This section presents the architecture of the Ocean Guard tool, designed to implement the requirements presented at section 2. Figure 1 showcases the components planned for Ocean Guard, following the reference architecture representation proposed by Ferreira et al. [6].
Figure 1: System Architecture of the Ocean Guard tool This architecture follows the same notation introduced by Ferreira et al. [6]. Rectangles represent applications or services, which execute continuously. Stacked rectangles represent pipelines, which execute a task on demand. Lastly, cylinders represent data storage, which may be databases of any type. Components are connected by arrows. Black arrows with a hollow tip illustrate the execution flow. They start and end in a component. Labeled arrows represent the trigger that starts a workflow, whereas unlabeled arrows represent the continuation of an existing workflow. Colored arrows with a filled tip illustrate the data flow. They appear in two types: solid arrows going to and from a data storage represent write and read operations, respectively; dotted arrows represent a sync or async request-response communication between components. Components are colored according to the data they produce: raw data  , source code  , executable artifacts  , ML-specific data  , ML models  , ML training metadata  , ML model predictions  , and ML model metrics  . Remaining standalone components  orchestrate the execution of others. Components are also grouped into subsystems, with their background colored according to the teams responsible for their development. Numbers  , roman numerals  and letters  are used as labels throughout section 3.
Data Collection.
To fulfill the requirements described in section 2, the Ocean Guard tool required comprehensive data acquisition capabilities. The Data Acquisition subsystem collects data from multiple heterogeneous sources: 3rd party Providers (1), Physical Sensors (2), and Data Crawlers (3). The data is stored in one of the two data stores: the Label Store (A) for known data about anomalies – which may come from the 3rd Party Providers (1) or the Web Application (1) – and theRaw Data Store (B) for unprocessed data.
@@
The next step in the implementation was to enable continuous training. Four components were developed for this purpose: a Synthetic Data Generation Pipeline (I); a Data Augmentation Pipeline (II); a Rule-Based Training Pipeline (III); and a ML-Based Training Pipeline (IV). The creation of these pipelines is an opportunity to introduce two critical model management tools: the Metadata Store (F) and the Model Registry (G), which together facilitate model lineage tracking and versioning.
@@
Currently, software engineers, data scientists, and machine learning engineers can access the Ocean Guard source code maintained in the centralized Code Repository (C). Developers can easily modify the source code by employing one of the pre-configured tools: Integrated Development Environments (IDEs) (4) for structured software engineering tasks, and Notebooks (5) for interactive prototyping and analysis. This approach accommodates the diverse workflows and preferences of a multidisciplinary team.
Crontinuous Delivery.
Following modern software engineering practices, a Continuous Integration (CI) Pipeline (v) is used to build and test code. After that, dedicated Continuous Delivery (CD) Pipelines (VI) deploy traditional software-based components, whereas specialized Continuous Delivery of Machine Learning (CD4ML) Pipelines (VII) do the same for ML-based components. These pipelines interact with a centralized Artifact Registry (D) that maintains versioned software components, containerized applications and serialized models, ensuring reproducibility and traceability throughout the development cycle.
@@ 4 API Architecture
This section dives deeper into the architecture of the Ocean Guard API, which corresponds to the API Prediction Service (8) in the architecture presented in section 3. The Ocean Guard API is the center of the ML-Enabled System, since it applies the rule-based and machine-learning based anomaly detection models over the data received by the system.
Figure 2: Component Architecture of the Ocean Guard API. It follows the hexagonal architecture [12]. The dark blue hexagon represents the Business Logic, where Domain Driven Design patterns such as Entities, Value Objects and Services [4] are implemented. The light blue hexagon represents the Dependencies, where the Ports and Adapters pattern is implemented as an Anti-Corruption Layer against external dependencies [12]. Rectangles illustrate the external dependencies themselves, with an orange rectangle representing a Reading Dependency and and yellow rectangle representing an Output Dependency. Lastly, circles represent the different Ports and Adapters themselves, one per dependency. Solid arrows, going from a port to an adapter, represent the inside-out dependency flow, going from the Core to Ports then Adapters.
Figure 2 represents the key components of the API according to a hexagonal architecture, also known as onion or clean architecture [12]. This architectural pattern is designed to enhance flexibility and maintainability of a system by ensuring that its core business logic remains independent of external technologies.
According to the hexagonal architecture pattern, the system can be divided into three key elements: core, ports and adapters. The system follows a dependency rule: As one moves inward towards the core, the level of abstraction and the scope of policy increases. Specifically, the deeper layers of the architecture encapsulate higher-level policies and overarching business rules, while the outer layers handle more concrete implementation details. The core, being the innermost layer, represents the most general and highest level of abstraction [12].
@@
Responsible for implementing the business logic. The Ocean Guard API follows the patterns of Domain Driven Design (DDD) [4]: entities and value objects represent the key objects related to the maritime domain. The services, also known as use cases [12], orchestrate the lifecycle of these objects.
@@
Responsible for establishing the contract and managing the communication between the core and the adapters. Within the Ocean Guard API, key components such as database repositories, dependency injection, security mechanisms, and web routers are defined as ports, facilitating the necessary integrations and ensuring that external concerns do not directly impact the core business logic.
@@ 5 Team Configuration and Development Workflow
As a ML-enabled system, developing the Ocean Guard tool requires multiple specialists, including software engineers, data scientists, and machine learning engineers [10]. Table 1 summarizes the project organizational structure, which consists of four teams working in parallel, yet collaborative streams.
Table 1: Ocean Guard Teams. The Research and Innovation teams focus on different AI-based and algorithmic techniques. Whilst the work from them is research- and innovation-oriented, their work directly contributes to the Ocean Guard API Model Adapter. The Core Development team focuses on the API whereas the UI Development focuses on the WebApp.
The project workflow is deliberately structured to minimize interdependency between teams. Each team operates within clearly defined boundaries whose goal is to allow parallel development without blocking dependencies. The integration points happen via the CI/CD pipeline [5], where independently developed components can be tested against predefined contracts before merging into the main development branch [16]. This design technique is further explored in section 6.
@@
It consists of Master students completing their theses. This team primarily focuses on exploring state-of-the-art techniques in the maritime domain. Although primarily research-oriented, their work directly contributes data products to the API, particularly to the API Model Adapter.
@@
It consists of technical trainees focused on the exploration and implementation of state-of-the-practice techniques. This team employs rapid prototyping methodologies within constrained timelines to deliver demonstrable improvements to the API Model Adapter.
@@
It consists of the authors of this paper focused on the fundamental aspects of the project infrastructure, including back-end development and infrastructure management. This team focuses on the API, databases, and model repository, with particular emphasis on refining the central API that serves as the integration point for all microservices.
@@
It consist of hired web developers specializing in user interface design and front-end development. This team focuses primarily on the WebApp, which will be used by the end users for maritime tracking activities.
@@ 6 Contract-based development
As an ML-enabled system with a microservice architecture, the Ocean Guard tool was a complex system, as discussed in section 3. A recommended modeling technique for systems with such an architecture is Design by Contract [7]: an approach that establishes clear interfaces and agreements between software components or teams, defining how they interact without requiring knowledge of each other’s internal implementation [7].
In the development of the Ocean Guard tool, contract-based development was systematically implemented across the boundaries of interaction of the four teams introduced at section 5. These contracts served as formal specifications defining interaction patterns without constraining the internal implementations [13].
@@ 6.1 Code Contracts
For this project, Code Contracts document the expected behavior between two services – owned by different teams – which interacting synchronously or asynchronously via the network, usually under the HTTP protocol. As such, these specifications include endpoints, request parameters, and expected response.
@@
  • Between Data Crawlers (3) and external data sources. Creating these contracts provides the Core Development team with an anti-corruption layer [4] against the external data sources. This allowing them to create tests that verify the expected behavior of 3rd party APIs, which may change without notice.
  • Between the API Prediction Service (8) and the Web Application (8). Creating this contract allows the UI Development team to mock the responses of the API Prediction Service (8). This allows them to develop the front-end in parallel with the back-end developed by the Core Development team.
@@ 6.3 Model Contracts
For this project, Model Contracts document the expected input and output of a rule-based or ML-based model, as well as the format on which a trained model should be stored. As such, these specifications include data types, file formats, and data scientist–defined hyperparameters.
There is one main model contract in the Ocean Guard tool, at the Model Store (B), between the Rule-Based Training Pipeline (III) and ML-Based Training Pipeline (IV) pipelines that create them, and both the Batch Prediction Pipeline (VIII) and API Prediction Service (9) that use them. These contracts ensure the Research and Innovation teams know what to create for the Core Development team to use in the Ocean Guard tool back-end.
@@ 7 Challenges
As an ML-enabled system, the Ocean Guard tool was intrinsically complex, as discussed in section 3. Any modification to one of its components – either on code, model, or data – can quickly cascaded through the system. Therefore, mitigating the drawbacks of coupling was an important architectural challenge, particularly since services were being implemented concurrently during development.
@@
As an ML-enabled system, the Ocean Guard tool required specialized developers, as discussed in section 5. Four teams worked simultaneously on different components, integrated via the Ocean Guard API. Therefore, keeping all team members oriented towards the same goal was an important organizational challenge, particularly due to the time-consuming nature of coordination activities.
@@
As an ML-enabled system with a microservices architecture, the Ocean Guard tool was a complex system, as discussed in section 3. Therefore, communicating the evolution of the Ocean Guard tool to internal and external stakeholders – with diverse technical backgrounds – was an important project management challenge, particularly since there were frequent feedback sessions due to its agile development process.
@@ 8 Lessons Learned
The challenges described in section 7 led to two lessons learned, related to applying techniques from the microservices literature that can be suitable for ML-enabled system development: contract-based design and ubiquitous language.
@@
As previously mentioned in section 6, the authors employed design by contract [7] during the development of the Ocean Guard tool, a recommended modeling technique for systems with a microservices architecture [16]. This approach facilitated the definition of clear and precise specifications between components, setting expectations for interfaces and behaviors.
By applying design by contract through the development process, it tackled two challenges above: coupling, by reducing potential integration issues and improving system cohesion, and alignment, making it easier for the team to discuss how to integrate the components of the Ocean Guard tool.
@@
As previously mentioned in section 3, the authors employed Domain-Driven Design (DDD) [4] during the development of the Ocean Guard tool, a recommended modeling technique for systems with a microservices architecture [14]. This approach facilitated the creation of a shared vocabulary between developers and stakeholders, grounding discussions about the evolution of the system.
By refining a ubiquitous language through the development process, it tackled two challenges above: communication, by deepening stakeholders’ and developers’ understanding of the system and improving delivery feedback, and alignment, making it easier for the team to propose how to evolve the components of the Ocean Guard tool.

### v2

@@ Abstract
This case study describes challenges and lessons learned on building Ocean Guard: a Machine Learning–Enabled System (MLES) for anomaly detection in the maritime domain. First, the paper presents the system’s specification and architecture. Ocean Guard was designed with a microservices architecture to enable multiple teams to work on the project in parallel. Then, the paper discusses how the developers adapted contract-based design to MLOps for achieving that goal. As an MLES, Ocean Guard employs code, model, and data contracts to establish guidelines between its services. This case study hopes to inspire software engineers, machine learning engineers, and data scientists to leverage similar approaches for their systems.
@@ 1 Introduction
Since 2015, the microservices architectural style has emerged as a promising approach to address challenges of modularity, scalability, and resilience by decomposing complex systems into smaller, independently deployable services with well-defined interfaces [13, 15]. This approach can be particularly suited to the maritime domain’s nature, where specialized components can be designed to handle different maritime data sources and analytical needs.
@@
This experience report describes the development of Ocean Guard: an ML-Enabled System (MLES) for anomaly detection in the maritime domain. Section˜2 details the functional and non-functional requirements considered for the system. Section˜3 gives an overview of the system’s microservices architecture. Section˜4 focuses on the Ocean Guard API, designed with a hexagonal architecture using Domain-Driven Design (DDD). Section˜5 shows the different teams involved in the development of the tool. Section˜6 describes how the developers adapted contract-based design to MLES for developing different components in parallel. Section˜7 lists the main challenges during the development of the Ocean Guard tool.Section˜8 then summarizes the key lessons learned, related to applying development techniques from microservices to MLES. Finally, section˜9 lists the main contributions of the paper and future work.
Unfortunately, given the sensitivity of the project, this paper does not address Ocean Guard data model or its anomaly detection techniques. Instead, it focuses on the high-level architecture design, which may inspire other researchers and practitioners developing similar ML-enabled systems.
@@ 2 System Specification
This section describes the specification of the Ocean Guard tool. Its goal is to analyze and detect anomalies across multiple types of data, unlike other commercially available tools. The section is divided into two parts: functional and non-functional requirements [14]. They were defined through interviews with the stakeholders of the project during the inception of the Ocean Guard tool.
@@ 2.1 Functional Requirements
Functional Requirements specify the observable behavior that a system should provide [17]. There are two actors related to the Ocean Guard tool. The following descriptions detail the main functionalities associated with them.
@@
    see geolocations (latitude, longitude, timestamp) of marine objects on a map,
@@ 2.2 Non-Functional Requirements
Non-Functional Requirements describe non-domain design considerations that influence the structure of a system and are important for its success [14]. There are four main non-functional requirements proposed for the Ocean Guard tool. The following descriptions detail the main characteristics the system should have to attend to its users.
@@
Compliance.
The Ocean Guard tool should be compliant with the latest European laws, including the General Data Protection Regulation (GDPR), and the Artificial Intelligence Act (AI Act). To make data and model management simpler to comply with different regulations, each stakeholder should be able to have their own installation of Ocean Guard.
@@ 3 System Architecture
This section presents the architecture of the Ocean Guard tool, designed to implement the requirements presented in section˜2. Figure˜1 showcases the components planned for Ocean Guard, following the reference architecture representation proposed by Ferreira et al. [5].
Figure 1: System Architecture of the Ocean Guard tool This architecture follows the same notation introduced by Ferreira et al. [5]. Rectangles represent either applications or services, which execute continuously. Stacked rectangles represent pipelines, which execute a task on demand. Lastly, cylinders represent data storage, which may be databases of any type. Components are connected by arrows. Black arrows with a hollow tip illustrate the execution flow. They start and end in a component. Labeled arrows represent the trigger that starts a workflow, whereas unlabeled arrows represent the continuation of an existing workflow. Colored arrows with a filled tip illustrate the data flow. They appear in two types: solid arrows going to and from a data storage represent write and read operations, respectively; dotted arrows represent a sync or async request-response communication between components. Components are colored according to the data they produce: raw data  , source code  , executable artifacts  , ML-specific data  , ML models  , ML training metadata  , and ML model predictions  . The remaining standalone components  orchestrate the execution of others. Components are also grouped into subsystems, with their background colored according to the teams responsible for their development. Numbers  , roman numerals  and letters  are used as labels throughout section˜3.
Data Acquisition.
To fulfill the requirements described in section˜2, the Ocean Guard tool required comprehensive data acquisition capabilities. This subsystem collects data from multiple heterogeneous sources: 3rd party Providers (1), Physical Sensors (2), and Data Crawlers (3). The data is stored in one of the two data stores: the Label Store (A) for known data about anomalies – which may come from the 3rd Party Providers (1) or the Web Application (9) – and the Raw Data Store (B) for unprocessed data.
@@
The next step in the implementation was to enable continuous training. Four components were developed for this purpose: a Synthetic Data Generation Pipeline (I); a Data Augmentation Pipeline (II); a Rule-Based Training Pipeline (III); and an ML-Based Training Pipeline (IV). The creation of these pipelines is an opportunity to introduce two critical model management tools: the Metadata Store (F) and the Model Registry (G), which together facilitate model lineage tracking and versioning.
@@
Currently, software engineers, data scientists, and machine learning engineers can access the Ocean Guard source code maintained in the centralized Code Repository (C). Developers can easily modify the source code by employing one of the pre-configured tools: Integrated Development Environments (IDEs) (4) for structured software engineering tasks, and Notebooks (5) for interactive prototyping and analysis. Providing standardized environments allows quick onboarding in the project and a seamless development experience.
Continuous Delivery.
Following modern software engineering practices, a Continuous Integration (CI) Pipeline (v) is used to build and test code. Thereafter, dedicated Continuous Delivery (CD) Pipelines (VI) deploy traditional software-based components, whereas specialized Continuous Delivery of Machine Learning (CD4ML) Pipelines (VII) do the same for ML-based components. These pipelines interact with a centralized Artifact Registry (D) that maintains versioned software components, containerized applications and serialized models, ensuring reproducibility and traceability throughout the development cycle.
@@ 4 API Architecture
This section dives deeper into the architecture of the Ocean Guard API, which corresponds to the API Prediction Service (8) in the architecture presented in section˜3. The Ocean Guard API is the center of the ML-Enabled System, since it applies the rule-based and machine-learning-based anomaly detection models over the data received by the system.
Figure 2: Component Architecture of the Ocean Guard API. It follows the hexagonal architecture [12]. The purple hexagon  represents the Business Logic, where Domain-Driven Design patterns such as Entities, Value Objects, Aggregates, and Services [4] are implemented. The blue hexagon  represents the Dependencies, where the Ports and Adapters pattern is implemented as an Anti-Corruption Layer against external dependencies [12]. Rectangles illustrate the external dependencies themselves, with a green rectangle  representing a Driving Dependency – which initiates actions in the API – while an orange rectangle  representing a Driven Dependency – whose actions are initiated by the API. Lastly, yellow circles  represent the Ports and red circles  represent the Adapters, one per dependency. Solid arrows, going from a port to an adapter represent the inside-out dependency flow, going from the Core to the Ports then to the Adapters.
Figure˜2 represents the key components of the API according to a hexagonal architecture, also known as onion or clean architecture [12]. This architectural pattern is designed to enhance flexibility and maintainability of a system by ensuring that its core business logic remains independent of external technologies.
According to the hexagonal architecture pattern, the system can be divided into three key elements: core, ports, and adapters. The system follows a dependency rule: As one moves inward towards the core, the level of abstraction and the scope of policy increases. Specifically, the deeper layers of the architecture encapsulate higher-level policies and overarching business rules, while the outer layers handle more concrete implementation details. The core, being the innermost layer, represents the most general and highest level of abstraction [12].
@@
Responsible for implementing the business logic. The Ocean Guard API follows the patterns of Domain-Driven Design (DDD) [4]: entities and value objects represent the key objects related to the maritime domain. The services, also known as use cases [12], orchestrate the lifecycle of these objects.
@@
Responsible for establishing the contract and managing the communication between the core and the adapters. Within the Ocean Guard API, key components such as database repositories, dependency injection, security mechanisms, and web routers are defined as ports, facilitating the necessary integrations and ensuring external concerns do not impact the core business logic.
@@ 5 Team Configuration and Development Workflow
As an ML-enabled system, developing the Ocean Guard tool requires multiple specialists, including software engineers, data scientists, and machine learning engineers [10]. Table˜1 summarizes the project organizational structure, which consists of four teams working in parallel, yet collaborative streams.
Table 1: Ocean Guard Teams. The Research and Innovation teams – composed of M.Sc. students and Eng.D. trainees – focus on different AI-based and algorithmic techniques. Whilst the work from them is research- and innovation-oriented, their work directly contributes to the Ocean Guard API Model Adapter. The Core Development team focuses on the API whereas the UI Development focuses on the WebApp.
The project workflow is deliberately structured to minimize interdependency between teams. Each team operates within clearly defined boundaries whose goal is to allow parallel development without blocking dependencies. The integration points happen via the CI/CD pipeline [7], where independently developed components can be tested against predefined contracts before merging into the main development branch [15]. This design technique is further explored in section˜6.
@@
It consists of Master of Science (M.Sc.) students completing their theses. This team primarily focuses on exploring state-of-the-art techniques in the maritime domain. Although research-oriented, their work directly contributes data products to the API, particularly to the API Model Adapter.
@@
It consists of Engineering Doctorate (Eng.D.) trainees focusing on the exploration and implementation of state-of-the-practice techniques. This team employs rapid prototyping methodologies within constrained timelines to deliver demonstrable improvements to the API Model Adapter.
@@
It consists of the authors of this paper focusing on the fundamental aspects of the project infrastructure, including back-end development and infrastructure management. This team focuses on the API, databases, and model repository, with particular emphasis on refining the central API that serves as the integration point for all microservices.
@@
It consists of hired web developers specializing in user interface design and front-end development. This team focuses primarily on the WebApp, which will be used by the end users for maritime tracking activities.
@@ 6 Contract-Based Development
Ocean Guard is a complex system, as discussed in section˜3. Design by Contract is a recommended modeling technique for systems with such an architecture [6]. It is an approach that establishes clear interfaces and agreements between software components or teams, defining how they interact without requiring knowledge of each other’s internal implementation [6].
In the development of the Ocean Guard tool, contract-based development was systematically implemented across the boundaries of interaction of the four teams introduced at section˜5. These contracts serve as a formal specification without constraining the internal implementations [15]. As the project evolves, contracts can be refined to keep teams aligned.
@@ 6.1 Code Contracts
For this project, code contracts document the expected behavior between two services – owned by different teams – that interact synchronously or asynchronously via the network, usually under the HTTP protocol. As such, these specifications include endpoints, request parameters, and expected responses.
@@
  • Between Data Crawlers (3) and external data sources. Creating these contracts provides the Core Development team with an anti-corruption layer [4] against the external data sources. This allows them to create tests that verify the expected behavior of 3rd-party APIs, which may change without notice.
  • Between the API Prediction Service (8) and the Web Application (8). Creating this contract allows the UI Development team to mock the responses of the API Prediction Service (8). This allows them to develop the front end in parallel with the back end developed by the Core Development team.
@@ 6.3 Model Contracts
For this project, model contracts document the expected input and output of a rule-based or ML-based model, as well as the format in which a trained model should be stored. As such, these specifications include data types, file formats, and data scientist–defined hyperparameters.
There is one main model contract in the Ocean Guard tool, at the Model Registry (G), between the Rule-Based Training Pipeline (III) and ML-Based Training Pipeline (IV) pipelines that create them, and both the Batch Prediction Pipeline (VIII) and API Prediction Service (9) that use them.
These contracts ensure the Research and Innovation teams know how to design their models for the Core Development team to use in the Ocean Guard tool back-end. This way, all teams know what the model may expect as input, what it should return as output, and which artifacts it may produce while executing. These contracts ensure seamless integration of the model during runtime.
@@ 7 Challenges
The Ocean Guard tool was intrinsically complex, as discussed in section˜3. Any modification to one of its components – either on the code, model, or data dimension – can quickly cascade through the system. Therefore, mitigating the drawbacks of coupling was an important architectural challenge, particularly since services were being implemented concurrently during development.
@@
The Ocean Guard tool required specialized developers, as discussed in section˜5. Four teams worked simultaneously on different components, integrated via the Ocean Guard API. Therefore, keeping all team members oriented towards the same goal was an important organizational challenge, particularly due to the time-consuming nature of coordination activities.
@@
The Ocean Guard tool had a diverse set of requirements, as discussed in section˜2. Therefore, communicating the evolution of the Ocean Guard tool to internal and external stakeholders – with diverse technical backgrounds – was an important project management challenge, particularly since there were frequent feedback sessions due to its agile development process.
@@ 8 Lessons Learned
The challenges described in section˜7 led to two lessons learned related to applying techniques from the microservices literature that can be suitable for ML-enabled system development: contract-based design and ubiquitous language.
@@
As previously mentioned in section˜6, this project employed design by contract [6] during the development of the Ocean Guard tool, a recommended modeling technique for systems with a microservices architecture [15]. This approach facilitated the definition of clear and precise specifications between components, setting expectations for interfaces and behaviors.
By applying design by contract through the development process, it tackled two challenges above: coupling, by reducing potential integration issues and improving system cohesion and alignment, making it easier for the team to discuss how to integrate the components of the Ocean Guard tool.
Applying design by contract for MLOps is a contribution of this paper. To develop Ocean Guard, this project took the idea of applying code contracts between microservices [15] to also use data and model contracts between components of MLES.
@@
As previously mentioned in section˜3, this project employed Domain-Driven Design (DDD) [4] during the development of the Ocean Guard tool, a recommended modeling technique for systems with a microservices architecture [13]. This approach facilitated the creation of a shared vocabulary between developers and stakeholders, grounding discussions about the evolution of the system.
By refining a ubiquitous language through the development process, it tackled two challenges above: communication, by deepening stakeholders’ and developers’ understanding of the system and improving delivery feedback and alignment, making it easier for the team to propose how to evolve the components of the Ocean Guard tool.
Creating a ubiquitous language for MLOps is a contribution of this paper. To develop Ocean Guard, this project took the idea of sharing common terms between bounded contexts [4] to also explain data- and model-related concepts for the stakeholders.